# Query-Guided Self-Supervised Summarization of Nursing Notes

### v1

@@ 3.3 Augmentation Blocks for the Context of the Patient
With these two augmentation blocks, the computation of the final decoder state $\mathbf{H}^{dec}\in\mathbb{R}^{(j+2)\times d}$ for generating the $(j+1)$th token in the summary of the note $N_{i}$ is abstracted as:
@@
$\mathbf{v}=\operatorname{LMH}(\mathbf{H}^{dec}),\hskip 14.22636pt\mathbf{v^{\prime}}=\operatorname{ST-GumbelSoftmax}(\mathbf{v}).$ (9)
4 Experiments
4.1 Data
We utilize MIMIC-III (Johnson et al., 2016), a widely used real-world EHRs database, for our experiments. In MIMIC-III, clinical notes in “NOTEEVENTS” table are organized by admission, and a single patient may have multiple admissions. Since the information in notes from different admissions of the same patient is discontinuous, we treat notes in each admission independently. We focus on nursing notes within the clinical notes. After the preprocessing, filtering and sampling (details in Appendix A.2), the number of nursing notes in the training, validation and test sets is 149015, 10001, 3079 and the corresponding numbers of admissions are 13893, 1000, 1156.
4.2 Types of Queries
This section presents queries used in our experiments, and more details can be found in Appendix A.3. Two principles are followed when determining the queries: (1) The query should be closely related to the patient and learnable by the query responder $\operatorname{R}$; (2) Data required to train $\operatorname{R}$ should be easily available without excessive data annotation. Below we propose four different types of queries. In each of these, the query responder $R$ is a classification model, which classifies patients according to a specific aspect of a patient’s status. Part of the training data is used to train the query responder $\operatorname{R}$. When using $\operatorname{R}$ to guide the summarization, we input the summary and the original note to predict the classification probabilities, and minimize the discrepancy between these predictions, as described in Section 3.2. As an additional query, we include a simple baseline by minimizing the semantic distance between the note and its summary, measured by cosine similarity.
Contrastive Next Note Prediction.
Given a nursing note pair $(N_{i},N^{\prime}_{i})$, we regard the query about whether $N^{\prime}_{i}$ is the successor note of $N_{i}$ as a prediction of the patient’s future status. To train the query responder R for the next note prediction, we create two note pairs for each nursing note, where the positive pair comprises the note and its successor in the sequence, and the negative pair contains the note and a randomly chosen non-consecutive note. If $N_{i}$ is the patient’s last nursing note, we use the patient’s discharge summary and a random note from other patients to construct the positive and negative pairs. The query is formulated as binary classification, and the output of $\operatorname{R}$ is the probability of each pair being the positive pair containing the consecutive notes.
Readmission Prediction.
Phenotype Classification.
Readmission Prediction and Phenotype Classification.
4.3 Experiment Settings
As the base model, we use BART-Large-CNN [^2] https://huggingface.co/facebook/bart-large-cnn, which is a BART model (Lewis et al., 2020) fine-tuned on CNN Daily Mail, specialized in text summarization. As the query responder, we use Clinical-Longformer (Li et al., 2022), chosen for its ability to handle a long context. It is fine-tuned on the selected queries. Hyperparameter settings for QGSumm and the query responder are presented in Appendix A.4.
Since our method is designed for scenarios where reference summaries are unavailable, we compare it with baselines in both zero-shot and few-shot settings: Zero-Shot: BART-Large-CNN (abbreviated as BART-zs), BioMistral-7B (Labrak et al., 2024) (abbreviated as BioMistral-zs), and GPT-4 (OpenAI, 2023); 10-Shot: BART-Large-CNN (abbreviated as BART-fs), Pegasus (Zhang et al., 2020a), and BioMistral-7B (abbreviated as BioMistral-fs). We use summaries generated by GPT-4 for the 10-shot fine-tuning and use one-shot in-context learning when prompting GPT-4 and BioMistral-7B. We also include results from two extractive methods, TextRank (Mihalcea and Tarau, 2004) and Lead-40%, for reference. In TextRank, we utilize MPNet (Song et al., 2020) to obtain sentence embeddings. In Lead-40%, we use the first 40% of the content of the note as a summary. More details about baselines and in-context learning/few-shot adaptation can be found in Appendix B.1.
4.4 Evaluation Metrics
Evaluating the quality of text summarization is challenging (Bhandari et al., 2020), especially when reference summaries are unavailable. Therefore, we employ multiple metrics covering different aspects of the summaries, providing a comprehensive evaluation.
Automatic Evaluation Metrics.
Metrics in the automatic evaluation are divided into three categories: 1) predictiveness, 2) factuality and consistency, and 3) conciseness.
Metrics for predictiveness assess whether the summary adequately contains patient key information, quantified as the ability to predict the patient’s condition using the summary as input. Specifically, we conduct readmission prediction and phenotype classification using summaries from baselines and our method. We employ the summaries generated by different methods to fine-tune the query responder, resulting in multiple predictors, one for each method. For readmission prediction, we report the weighted F1 and F1 of the positive class (“being readmitted”), and for phenotype classification, we report the F1-Macro score.
For consistency and factuality, we consider: (1) UMLS-Recall. This metric measures the biomedical information consistency by comparing the set of medical concepts in the summary with that in the original note. We employ QuickUMLS (Soldaini and Goharian, 2016) to extract Unified Medical Language System (UMLS) biomedical concepts from the nursing note and its summary. Recall is the proportion of concepts in the original note that are present in the summary. (2) UMLS-FDR. FDR denotes False Discovery Rate. Analogously to UMLS-Recall, we compute the proportion of medical concepts in the summary that do not appear in the original note. (3) FactKB. With FactKB (Feng et al., 2023), an evaluation model measuring the factuality of a summary and its original text, we evaluate whether the summaries are consistent with the nursing notes from the perspective of their overall semantic information. (4) BARTScore. This metric evaluates the general consistency of summaries in a text-generation manner using BART, which also considers aspects such as the structure, coherence, and fluency of the summary (Yuan et al., 2021).
Finally, we report the length of the generated summary as a percentage of the original note’s length to assess conciseness. We do not enforce a strict maximum length for baselines because we believe the model should be capable of determining the appropriate length autonomously.
Metrics used in the manual evaluation by a clinician
Without a reference summary, automatic evaluation metrics may not fully capture the quality of the summary. Therefore, we invite a clinician to conduct manual evaluation of the summaries of 25 nursing notes. The clinician evaluates summaries from multiple methods in a blinded and randomized order. Each summary is rated on four aspects: (1) Informativeness: Whether the summary adequately captures essential information regarding the patient’s condition in the original note; (2) Fluency: Whether the summary is well-written and easy to understand. (3) Consistency: How well the summary aligns with the original nursing note in factuality. (4) Relevance: It evaluates the conciseness of the summary and whether it contains unnecessary information. The score for each aspect ranges from 1 to 5. More detailed grading criteria are presented in Appendix B.2.
@@ 5 Results and Discussion
{adjustwidth}
-1cm Type Method Predictiveness Consistency and Factuality Conciseness Readmission Phenotype UMLS-Recall UMLS-FDR BARTScore FactKB Length Weighted F1 F1 Macro F1 Orig. Notes 85.20.5 19.71.9 28.70.5 - - - - - Zero-Shot BART-zs 78.80.4 11.10.9 20.50.3 36.49.0 8.706.2 -1.890.31 0.780.16 31.9% GPT-4 85.60.6 21.52.0 23.60.6 59.28.3 44.27.6 -3.130.47 0.770.17 53.6% BioMistral-zs 80.10.6 10.71.3 21.40.4 55.49.9 50.08.7 -2.800.45 0.680.14 69.2% 10-Shot BART-fs 82.20.5 14.41.3 21.10.4 52.57.3 44.57.1 -2.720.36 0.760.15 65.0% BioMistral-fs 81.70.4 10.21.1 22.00.4 57.210.2 49.17.8 -2.970.43 0.700.15 68.8% Pegasus 80.50.8 12.51.8 18.30.6 35.18.4 52.67.7 -3.070.40 0.700.18 57.4% QGSumm -Similarity 79.50.6 12.01.2 22.40.4 53.17.2 20.76.7 -2.220.31 0.820.13 51.7% -NextNote 80.80.6 11.71.4 23.20.6 56.48.0 35.27.1 -2.320.33 0.770.11 49.3% -Readmission 82.40.5 18.21.6 23.90.5 58.27.5 22.76.5 -2.300.37 0.780.14 46.2% -Phenotype 81.90.5 13.41.5 25.60.6 58.57.4 36.26.9 -2.340.35 0.790.13 48.0% -Re+Ph 84.20.5 17.21.6 25.10.5 58.87.9 24.16.4 -2.260.35 0.800.14 48.2% Extractive Lead-40% 83.10.6 12.61.5 21.70.5 42.76.7 0.302.6 -0.870.11 0.990.06 40.0% TextRank 81.90.7 14.41.7 23.30.5 58.57.9 0.081.4 -0.900.15 0.950.12 51.9%
Table 1: Results of automatic evaluation. Lower values are better for Length and UMLS-HR, higher values for the other metrics. The subscripts denote standard deviation. “Orig. Notes” means using original nursing notes as such for readmission and phenotype prediction. “Re+Ph” means using “Readmission Prediction and Phenotype Classification” as the query. Results from best and 2nd best method under each metric are bolded and underlined. Extractive methods are for reference and not considered in comparison.
5.1 Results of the Automatic Evaluation
@@
Results are shown in Table 1. In the readmission prediction task, GPT-4 performs best, producing summaries that enable more accurate prediction of a patient’s status. Our method also exhibits excellent performance, surpassing all few-shot methods. The main reference for our method is BART-zs, as it is the base model in our method, and hence represents performance without the proposed novel components. We see that our method outperforms BART-zs significantly in weighted F1 score (84.2 vs. 78.8) and F1 score of the positive class (18.2 vs. 11.1). This shows the effectiveness of the adaptation strategy guiding the model with useful queries. Interestingly, we find that using the summary from GPT-4 for this task outperforms using the original notes. Similarly, the summary from our method has performance close to that of using the original notes, highlighting the importance of high-quality summaries. In phenotype classification, our method with the query focusing on patients’ phenotype performs the best, outperforming BART-zs in Macro F1 (25.6 vs. 20.5). Even when using the similarity alone as a guiding signal, our method still is better than BART-zs (22.4 vs. 20.5) or BART-fs (22.4 vs. 21.1). Although specialized in text summarization, Pegasus has weak performance on all predictiveness metrics.
Conciseness, Consistency, and Factuality.
As shown in Table 1, there is an expected trade-off between UMLS-Recall and the summary’s length. Our method strikes a good balance between medical information consistency (measured by UMLS-Recall) and conciseness. GPT-4 captures more medical information, but achieves this with summaries which are less concise. Conversely, BART-zs can produce concise summaries but fails to adequately capture medical concepts. Even if the 10-shot fine-tuning clearly improves predictiveness and UMLS-Recall, BART-fs still struggles to generate a concise summary. Similar to BART-fs, both BioMistral-zs and BioMistral-fs tend to produce summaries that are not concise.
Summaries generated from BART-zs maintain high levels in factuality (measured by UMLS-HR and FactKB) and general consistency (measured by BARTScore). Our method also has strong performance on relevant metrics. We find that although LLMs, such as GPT-4 and BioMistral, excel in language understanding, they do not perform well on factuality and general consistency. One possible reason is their tendency to rephrase or even expand upon the original notes, potentially introducing inconsistent information. However, the metrics can be influenced by text style and layout, which may cause summaries that are more different from the original note to score relatively lower, even if they are more fluent. For this reason, our model also scores lower than the base model BART-zs on some metrics. The results and limitations will be further discussed in Sections 5.4 and 5.5.
Effectiveness of the Query Guidance.
According to the results shown in Table 1, the performance with different queries varies. We can observe: (1) Regarding predictiveness, employing queries closely related to patients and focusing on readmission and phenotype information yield superior performance compared to other queries. As expected, the method with phenotype-related queries performs the best in phenotype classification, while the method with readmission-related queries is the best in readmission prediction. This highlights the effectiveness of guiding the summarization with queries, and different queries enable the summary to concentrate on different aspects of the original note. (2) Using similarity as guidance can produce summaries that are more similar to the original notes, resulting in higher scores on general consistency and factuality. However, summaries generated under this configuration tend to be longer and often sacrifice predictiveness and informativeness regarding medical concepts, demonstrating the limitations of the unconstrained guidance signal. (3) When employing joint guidance with both readmission and phenotype information, our method consistently achieves excellent performance across all metrics. This indicates that combining different guidance signals can help in producing better summaries, and further research is needed to explore this aspect in depth.
5.2 Results of the Manual Evaluation by a Clinician
@@
To avoid excessive manual work, we select three baselines to include in the manual evaluation: BART-zs, GPT-4, and BioMistral-fs. The justification for selecting these methods is two-fold: First, BART-zs is the base model in our method and hence the main baseline, demonstrating the benefits of the novel components. Second, GPT-4 and BioMistral-fs are well-known strong baselines and they had good performance in the automatic evaluation. We use “Readmission Prediction and Phenotype Classification” as the query for our method. Average scores for each method across four metrics are shown in Figure 4.
QGSumm vs BART-zs.
@@
Our method significantly outperforms the base model BART-zs on all four metrics. This indicates that the proposed domain adaptation strategy enables the model to generate higher-quality summaries from the medical personnel’s perspective, containing refined and important patient information with fewer hallucinations and increased readability. Although on average the summaries generated by our method are longer than those produced by the base model, our method achieves a higher relevance score from the clinician, suggesting the base model struggles to identify key information and focuses on unnecessary details. Our model can effectively enhance this aspect.
QGSumm vs GPT-4 and BioMistral-fs.
@@
Due to the LLMs’ strong language understanding capability and sufficient medical knowledge, GPT-4 and BioMistral-fs can adequately summarize key information in nursing notes, receiving approximately the same average score in Informativeness as our method. Additionally, they excel in generating fluent summaries by rephrasing and clarifying abbreviations, receiving a slightly higher Fluency from the clinician than our method. However, the rephrasing can introduce factual inconsistencies, and the tendency to infer additional content may reduce factuality. Consequently, our model has a higher average score in Consistency, which is essential in the clinical setting. Furthermore, it generates more concise summaries, leading to a higher average score in Relevance. However, due to the small sample size, the only statistically significant difference in these comparisons was the improvement of our method compared to Biomistral-fs in consistency, and further work is needed for conclusive results.
Calculating the Significance.
We calculated the significance in Figure 4 using a two-tailed Binomial test on the pairwise win-rates. In detail, we count the number of nursing notes where our method has a score higher or lower than a comparison method and test for the null hypothesis that the win-probability is 0.5.
Figure 4: Results of the manual evaluation by a clinician. The average scores in four metrics across 25 summarized notes are reported for QGSumm, BART-zs, GPT-4, and BioMistral-fs. “*” denotes the result of the significance test.
5.3 Effectiveness of Augmentation Blocks
We analyze the effects of the proposed augmentation blocks through an ablation study. We consider three settings: removing the Patient Information Augmentation block (denoted as w/o PIA); removing the Temporal Information Fusion block (denoted as w/o TIF); removing both blocks (denoted as w/o PIA+TIF). We use “Readmission Prediction and Phenotype Classification” as the query in our method. The results of the ablation study are shown in Table 2 The decreased weighted F1 and macro F1 scores indicate that both augmentation blocks enhance the predictiveness of summaries. This implies that information in patient metadata and previous notes can effectively prompt the inference of the current and future status of patients. Removing the TIF block causes a larger decrease in the F1 scores, suggesting that temporal information is more important than the patient’s metadata in guiding the generation of summaries to focus on the progression of the patient’s status.
On the other hand, the incorporation of patient metadata can lead to more faithful summaries, as the removal of PIA degrades more the performance on UMLS-FDR and FactKB, which are related to factuality. In contrast, the TIF block does not have a significant impact on the factuality. However, according to the UMLS-Recall score, it encourages the model to capture more medical information, thereby improving the consistency of the summary.
Table 2: Results of the ablation study. We present scores of five metrics for QGSumm, and show the change in the value of the metric after removing different augmentation blocks. $\downarrow$ denotes a decrease in the score and $\uparrow$ denotes an increase. We see that the augmentations are consistently useful.
@@ 5.3 Effectiveness of Augmentation Blocks
Weighted F1 Macro F1 UMLS-Recall UMLS-FDR FactKB
QGSumm 84.2 25.1 58.8 24.1 0.80
w/o PIA $\downarrow 2.6$ $\downarrow 1.4$ $\downarrow 1.4$ $\uparrow 4.7$ $\downarrow 0.04$
w/o TIF $\downarrow 4.1$ $\downarrow 1.6$ $\downarrow 3.8$ $\uparrow 1.9$ $\downarrow 0.01$
w/o PIA+TIF $\downarrow 4.8$ $\downarrow 2.3$ $\downarrow 4.4$ $\uparrow 5.6$ $\downarrow 0.04$
5.4 Case Study
Figure 5: One artificial nursing note and its summaries from QGSumm, BART-zs, GPT-4, and BioMistral-fs, respectively.
One artificial nursing note and its corresponding summaries generated by QGSumm, BART-zs, GPT-4, and BioMistral-fs are presented in Figure 5. In the original nursing note, the content highlighted in blue indicates information included in the summary generated by our approach. We can see that our approach captures most of the important patient information. However, some details, such as cardiovascular and respiratory conditions, are overlooked. The summary from BART-zs only covers information from the first half of the nursing note, suggesting the limitation in understanding long context. Summaries from GPT-4 and BioMistral-fs contain more patient information but lack conciseness. These models achieve fluency by rephrasing notes and expanding abbreviations. However, BioMistral-fs struggles with maintaining factuality, often excessively reasoning about the patient’s personal information and condition.
5.5 Discussion

### v2

@@ 3.3 Augmentation Blocks for the Context of the Patient
With these two augmentation blocks, we can obtain the final decoder state $\mathbf{H}^{dec}\in\mathbb{R}^{(j+2)\times d}$, which is input to the language modeling head for generating the $(j+1)$th token in the summary. The computation is shown in A.5
4 Experimental setup
4.1 Data
We use MIMIC-III (Johnson et al., 2016), an EHR database with clinical notes organized by admission. Notes are treated independently per admission due to the discontinuity between admissions of the same patient. We focus on the nursing notes within the clinical notes. After preprocessing (details in A.2), the numbers of notes in the training, validation and test sets are 149015, 10001, 3079 and the corresponding numbers of admissions are 13893, 1000, 1156.
4.2 Types of Queries
Two principles guide the query selection: (a) The query should be related to the patient and learnable by the query responder $\operatorname{R}$; (b) Training data for $\operatorname{R}$ should be easily available. We propose two queries: Readmission Prediction, predicting if a patient will be readmitted within 30 days post-discharge, and Phenotype Classification, identifying which of 25 phenotypes the patient has (Harutyunyan et al., 2019). In the experiments, we combine these queries.
The query responder $\operatorname{R}$ is a classifier that predicts an answer to the query. Part of the training data is used to train $\operatorname{R}$. When using $\operatorname{R}$, we input the summary or the original note to predict the classification probabilities, and minimize the discrepancy between these predictions (see Section 3.2). As a baseline we include a query by minimizing the cosine similarity between the note and its summary. More details on queries and their combination are in A.3.
4.3 Detailed Settings
We use BART-Large-CNN [^1] https://huggingface.co/facebook/bart-large-cnn as the base model, a BART model (Lewis et al., 2020) fine-tuned for text summarization. For the query responder, we use Clinical-Longformer (Li et al., 2022), as it can handle long contexts, and fine-tune it on the selected queries. Hyperparameters for the models are specified in A.4.
Since our method is designed for scenarios where reference summaries are unavailable, we compare our method with baselines in zero-shot settings: BART-Large-CNN (BART) and Pegasus (Zhang et al., 2020a); and in 1-shot in-context learning prompting using GPT-4 (OpenAI, 2023) and BioMistral-7B (Labrak et al., 2024) (BioMistral). Consistent with QGSumm, we fine-tune the encoder of BART and Pegasus to reconstruct notes, enhancing their performance. We consider BART as the most relevant baseline for our method as it is the base model in our method, and hence represents performance without the proposed novel components.
For a fair comparison, we use six different prompts when evaluating GPT-4 and BioMistral: (1) the original prompt without additional information; (2) with patient information; (3) indicating the summary is for readmission prediction; (4) for phenotype classification; (5) for both readmission prediction and phenotype classification; and (6) including all previous notes as temporal information. The original prompt is shown in the main results due to negligible performance differences between prompts (1)-(5), and poorer performance with the prompt (6). A single summary example is included in the prompt to ensure that the generated summaries align with structural requirements. The content of prompts and full results for five different prompts are in C.5. Additional results with few-shot fine-tuning settings and extractive methods are in C.4, and more details about baselines in B.1.
4.4 Evaluation Metrics
Evaluating the quality of text summarization is challenging (Bhandari et al., 2020), especially in specialized fields and without reference summaries. Therefore, we employ multiple metrics to provide a comprehensive evaluation.
Automatic Evaluation Metrics.
Metrics in the automatic evaluation are divided into three categories: 1) factuality and consistency, 2) predictiveness, and 3) conciseness.
For consistency and factuality, we consider: (1) UMLS-Recall. We use QuickUMLS (Soldaini and Goharian, 2016) to extract Unified Medical Language System (UMLS) biomedical concepts from the nursing note and its summary. Recall is the proportion of concepts in the original note that are present in the summary. (2) UMLS-FDR. Denotes False Discovery Rate, quantifying the proportion of medical concepts in the summary that are not present in the original note. (3) FactKB. Evaluates factual consistency of summaries based on overall semantic information (Feng et al., 2023).
Predictiveness metric assesses whether the summary adequately contains patient key information, quantified as the ability to predict the patient’s condition using the summary. Specifically, we conduct readmission prediction and phenotype classification using summaries from baselines and QGSumm. For a fair comparison, we use summaries from each method to fine-tune its respective classifier for the down-stream task (similar to the query responder). This ensures the best evaluation result on predictiveness for each method, as illustrated in B.3. For readmission prediction, we report the weighted F1 and F1 of the positive class (“being readmitted”), and for phenotype classification, we report the F1-Macro.
Finally, we report the summary length as a percentage of the original note’s length to assess conciseness. We do not enforce a strict maximum length for baselines as the model should be capable of determining the appropriate length autonomously.
Metrics used in the manual evaluation by a clinician.
Without a reference summary, automatic evaluation metrics may not fully capture the quality of the summary. Therefore, we invite a clinician to manually evaluate the summaries generated from 25 nursing notes by 4 summarization method (25$\times$4 summaries in total). The summaries are shown to the clinician in a randomized order and without showing the method that created them. Each summary is rated on four aspects from 1 to 5: (1) Informativeness: Whether the summary adequately captures essential information regarding the patient’s condition; (2) Fluency: Whether the summary is well-written and easy to understand. (3) Consistency: How well the summary aligns with the nursing note in factuality. (4) Relevance: Whether the summary is concise and not contain unnecessary information. Detailed grading criteria are in B.2.
{adjustwidth}
@@ 4.4 Evaluation Metrics
Table 1: Results of automatic evaluation. Lower values are better for Length and UMLS-FDR, higher values for the other metrics. The subscripts denote standard deviation. “Orig. Notes” means using original nursing notes as such for readmission and phenotype prediction. “Re+Ph” means using “Readmission Prediction and Phenotype Classification” as the query. Results from best and 2nd best method under each metric are bolded and underlined.
Type Method Consistency and Factuality Conciseness Predictiveness
UMLS-Recall UMLS-FDR FactKB Length Readmission Phenotype
Weighted F1 F1 Macro F1
Orig. Notes - - - - 85.20.5 19.71.9 28.70.5
Baselines BART 36.49.0 8.706.2 0.780.16 31.9% 78.80.4 11.10.9 20.50.3
GPT-4 59.28.3 44.27.6 0.770.17 53.6% 85.60.6 21.52.0 23.60.6
BioMistral 55.49.9 50.08.7 0.680.14 69.2% 80.10.6 10.71.3 21.40.4
Pegasus 32.28.7 47.47.9 0.650.12 51.9% 77.10.4 9.30.8 17.50.5
QGSumm -Similarity 53.17.2 20.76.7 0.820.13 51.7% 79.50.6 12.01.2 22.40.4
-Re+Ph 58.87.9 24.16.4 0.800.14 48.2% 84.20.5 17.21.6 25.10.5
[…]
fig:avg_score Refer to caption
Figure 4: Results of the manual evaluation by a clinician. The average scores in four metrics are reported for QGSumm, BART, GPT-4, and BioMistral. “*” denotes the result of the significance test, calculated using a two-tailed Binomial test on the pairwise win-rates, i.e., we count the number of notes where QGSumm has a score higher or lower than a comparison method and test for the null hypothesis that the win-rate is 0.5.
@@ 5 Results and Discussion
5.1 Automatic evaluation
Conciseness, Consistency, and Factuality.
There is a trade-off between UMLS-Recall and summary length (conciseness). As shown in Table LABEL:tab:result1, our method balance medical information consistency (measured by UMLS-Recall) and conciseness. GPT-4 captures more medical information but is less concise, whereas BART produces concise but less informative summaries. Similar to GPT-4, BioMistral tends to produce summaries that are not concise. Summaries from BART maintain high levels in factuality (UMLS-FDR and FactKB). Our method also has strong performance on relevant metrics. We find that although LLMs, such as GPT-4 and BioMistral, excel in language understanding, they do not perform well on factuality. One possible reason is their tendency to rephrase or expand the original notes, potentially introducing inconsistent information, i.e. hallucination, which can cause a relatively high UMLS-FDR.
@@ 5.1 Automatic evaluation
Results are shown in Table LABEL:tab:result1. In the readmission prediction task, our method closely follows GPT-4, surpassing all other baselines. We see that our method outperforms BART significantly in weighted F1 score (84.2 vs. 78.8) and F1 score of the positive class (17.2 vs. 11.1). This shows the effectiveness of the adaptation strategy guiding the model with useful queries. Interestingly, we find that using the summary from GPT-4 for this task outperforms using the original notes, and our method’s summaries perform nearly as well, highlighting their quality. In phenotype classification, our method performs the best, outperforming BART in Macro F1 (25.1 vs. 20.5). Even when using similarity alone, it surpasses BART (22.4 vs. 20.5). Notably, our method’s training objective does not involve true labels of the queries, and the query responder remains frozen during training, preventing query data leakage into $\operatorname{M}$. Instead, the strong predictiveness of our summaries follows from emphasizing relevant patient-related information present in the original notes. Although specialized in text summarization, Pegasus has weak performance on all predictiveness metrics.
5.2 Manual Evaluation by a Clinician
@@
To avoid excessive manual work, we select 3 baselines: BART, GPT-4, and BioMistral in addition to QGSumm. BART was selected as the main baseline, as it is the base model in our method but without the novel components. GPT-4 and BioMistral are selected due to their strong performance in the automatic evaluation. Average scores for each method across four metrics are shown in Figure LABEL:fig:avg_score. A case study comparing example summaries from these methods can be found in C.1.
QGSumm vs BART.
@@
QGSumm significantly outperforms BART on all four metrics. This indicates that our domain adaptation strategy enables the model to generate higher-quality summaries from the medical personnel’s perspective, containing refined and important patient information with fewer hallucinations and increased readability. Despite producing longer summaries, our method achieves a higher relevance score from the clinician, suggesting the base model struggles to identify key information and focuses on unnecessary details. Our model can effectively enhance this aspect.
QGSumm vs GPT-4 and BioMistral.
@@
GPT-4 and BioMistral perform similarly to QGSumm in Informativeness but excel slightly in Fluency by rephrasing and clarifying abbreviations. However, the rephrasing can introduce factual inconsistencies, and the tendency to infer additional content may reduce factuality. An example of this is provided in the case study in C.1. Consequently, QGSumm scores higher in Consistency, which is essential in the clinical setting. Furthermore, it generates more concise summaries with higher scores in Relevance. However, due to the small sample size, the only statistically significant difference in these comparisons was the improvement of QGSumm compared to Biomistral in consistency, and further work is needed for more conclusive results.
[…]
@@
tab:ablation_study Weighted F1 Macro F1 UMLS-Recall UMLS-FDR FactKB QGSumm(-Re+Ph) 84.2 25.1 58.8 24.1 0.80 w/o PIA $\downarrow 2.6$ $\downarrow 1.4$ $\downarrow 1.4$ $\uparrow 4.7$ $\downarrow 0.04$ w/o TIF $\downarrow 4.1$ $\downarrow 1.6$ $\downarrow 3.8$ $\uparrow 1.9$ $\downarrow 0.01$ w/o PIA+TIF $\downarrow 4.8$ $\downarrow 2.3$ $\downarrow 4.4$ $\uparrow 5.6$ $\downarrow 0.04$Readmission $\downarrow 1.8$ $\downarrow 1.2$ $\downarrow 0.6$ $\downarrow 1.4$ $\downarrow 0.02$Phenotype $\downarrow 2.3$ $\uparrow 0.5$ $\downarrow 0.3$ $\uparrow 12$ $\downarrow 0.01$
Table 2: Results of the ablation study. We show the change in the value of the metric after removing different augmentation blocks/using different queries. $\downarrow$ denotes a decrease in the score and $\uparrow$ denotes an increase.
5.3 Ablation Study
We conducted an ablation study to: (1) evaluate the performance of the model without the proposed augmentation blocks: without the Patient Information Augmentation (w/o PIA), without the Temporal Information Fusion (w/o TIF), and without both blocks (w/o PIA+TIF); and (2) evaluate the effectiveness of the query guidance by using “Readmission Prediction” and “Phenotype Classification” as separate queries, instead of the combined query used earlier.
Table LABEL:tab:ablation_study shows that the removal of augmentation blocks causes weighted F1 and macro F1 scores decrease in all settings, indicating that both blocks enhance the predictiveness of the summary. Removing TIF results in a larger score drop in F1 scores, underlining the importance of temporal information to understand the patient’s current status. Conversely, the removal of PIA degrades the performance on factuality (UMLS-FDR and FactKB) more than the removal of TIF, while the removal of TIF affects more the UMLS-Recall, i.e., fewer medical concepts are captured. On the other hand, employing different queries allows the model to focus on different aspects of the original note, resulting in summaries varying across different metrics. While there are some exceptions, the general trend is that the combined query yields overall better performance than using either query alone. Full results and a more detailed analysis are presented in C.6.
5.4 Discussion and Conclusion